# ZeroFlow: Overcoming Catastrophic Forgetting is Easier than You Think

### v3

@@ 4 ZeroFlow Benchmark
This section delves into the empirical performance of ZO optimization in overcoming catastrophic forgetting. Our ZeroFlow includes evaluating average performance along incremental stages, last stage accuracy, forgetting, and efficiency, considering the complexity of different datasets, and considering various model types.
4.1 Benchmark Setups
@@
Forgetting scenarios, schemes, and models. We assigned the evaluation to a common catastrophic forgetting scenario, which is continual learning. We focus on four fine-tuning tasks, considering their complexity from the typical dataset CIFAR and CUB to dataset ImageNet-A, and OmniBenchmark, which has a large domain gap. For the method used on these tasks, we explore two SOTA schemes: EASE and APER. Both models are initially configured with ViT-B/16-IN21K and subsequently fine-tuned on four datasets. Following (Zhou et al., 2023a), each dataset is split equally into 10 tasks by class. For example, OmniBenchmark contains 300 classes, with 30 classes learned in each incremental stage. No memory is allowed for storing past examples.
@@
Benchmark setup and details. To evaluate the application of ZeroFlow in forgetting scenarios, we include the methods described in Section 3.1, specifically ZO (Ghadimi & Lan, 2013; Malladi et al., 2023), Sign (Liu et al., 2019), and Conserve (Kim et al., 2021; Zhang et al., 2024), in comparison with their FO counterparts using SGD and Adam optimizers (Chen et al., 2019). Additionally, as highlighted in (Zhang et al., 2024), Forward-Grad (Ren et al., 2022) which relies on forward mode automatic differentiation, potentially becomes a missing but competitive forward pass baseline. In a nutshell, ZeroFlow covers 7 forward pass methods: ZO-SGD, ZO-SGD-Sign, ZO-SGD-Conserve, ZO-Adam, ZO-Adam-Sign, ZO-Adam-Conserve, Forward-Grad.
@@
Evaluation metrics. Overall, we use two sets of metrics as ZeroFlow evaluation measures: accuracy and efficiency. Accuracy measures include average accuracy over all tasks, the last-task accuracy, and a forgetting metric. Efficiency measures include memory efficiency (GPU usage), query efficiency, and runtime efficiency. Together, these metrics provide insights into the resources required for ZO optimizations to overcome forgetting.
@@
4.2 Evaluation Results of ZeroFlow
ZeroFlow evaluation on continual learning. In Table 1, we evaluate the performance of different BP-free and BP-based (FO-SGD and FO-Adam) methods in a typical forgetting scenario (continual learning). We use two SOTA models as examples (EASE (Zhou et al., 2024b) and APER (Zhou et al., 2023a)) and investigate SGD and Adam optimizers, seven forward pass methods, and four commonly used datasets. Several observations are listed below,
@@ 4.2 Evaluation Results of ZeroFlow
ZeroFlow helps manage memory and runtime. In Table 2, we compare the efficiency of various ZO and FO optimizers in overcoming catastrophic forgetting, focusing on two main aspects: memory cost (in GB) and runtime cost (in seconds). First, naïve ZO optimization reduces memory cost by approximately fivefold compared to FO optimization. Meanwhile, ZO optimization decreases runtime cost per iteration by around 50% compared to FO, which significantly broadens the applicability of ZO optimization in overcoming forgetting. Second, ZO and sign methods show similar levels of efficiency in terms of memory efficiency and runtime costs. Although the number of queries can impact the efficiency of ZO methods, it does not destroy memory management efficiency. Third, Conserve also features efficient memory management, though its runtime cost is about 2x longer than the ZO. This may explain why Conserve performs better in a few cases, as shown in Table 1.
@@
Figure 6: The Trajectory of Different Optimization during Overcoming Forgetting. Red denotes the minimum of new task, blue denotes the minimum of old task. The trajectory taken when using the total loss from both tasks (cyan).
Trade-off between performance and query number. As shown in Figure 5, we investigate the impact of query numbers on optimization performance, comparing SGD and Adam optimizers in the zeroth-order setting. Both optimizers demonstrate improved performance as query numbers increase across [1,2,4,8,16,32], suggesting that additional function evaluations enable more accurate gradient estimation. The results suggest that in scenarios where function evaluation costs are manageable, higher query numbers can yield substantially better performance, with Adam being particularly effective at leveraging the additional gradient information for enhanced optimization outcomes.
5 Insights and Discussions
As shown in Figure 6, we visualized the optimization trajectories of both forward pass and backpropagation methods. Our analysis reveals several key insights:
Distinct trajectory characteristics of FO and ZO. FO approaches (Figure 6, 6) show smoother optimization paths due to their access to exact gradient information. In contrast, ZO methods demonstrate varying degrees of exploration behavior through trajectory jitter. This exploration pattern is particularly pronounced in ZO-Adam variants compared to ZO-SGD variants, indicating that the base optimizer choice significantly influences the exploration-exploitation trade-off during optimization.
Path characteristics in ZO optimization. Comparing base ZO methods with their $q=4$ counterparts (Figure 6 vs 6, Figure 6 vs 6), we observe that increasing query numbers leads to smoother trajectories, suggesting that more queries help provide more stable gradient estimates. The Sign variants (Figure 6, 6) demonstrate more pronounced oscillations in their trajectories, particularly visible in the ZO-Adam-Sign case. In contrast, the conservative variants (Figure 6, 6) maintain relatively stable paths that better balance between the old and new task minima.
6 New Enhancement to Mitigate Forgetting
Figure 7: Effectiveness of Hybrid ZO in Overcoming Forgetting. In Hybrid ZO, backward benefits from forward pass.
(a) FO-SGD
(b) ZO-SGD
Table 3: Effectiveness of Historical Estimation in Mitigating Forgetting. Proportion of 0% denotes that the plain optimizer ZO-SGD.
Metrics Proportion
0% 20% 40% 60% 80%
Avg 57.87 58.90 58.76 58.34 57.83
Last 48.32 49.04 48.84 48.42 48.10
Fgt 11.08 11.79 11.78 11.60 11.57
Table 4: Effectiveness of Sparsity-induced Estimation in Overcoming Forgetting. Proportion of 0% denotes the plain ZO-SGD.
Ratio 0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Avg 57.87 59.17 59.46 59.29 59.39 59.45 59.26 59.39 59.38 59.47
Last 48.32 48.58 49.05 48.72 48.91 49.24 49.11 49.05 49.11 49.24
Fgt 11.08 12.65 12.17 12.76 12.53 12.37 12.36 12.54 12.46 12.33
7 Conclusion
This paper introduces ZeroFlow, a benchmark study that probes a series of forward pass methods for overcoming catastrophic forgetting. This work resorts to an easier way (no need for backpropagation and activation storage) to overcome forgetting. Concretely, our benchmarks include various forward pass methods, forgetting scenarios, and evaluation metrics. We also reveal the overlooked optimization principles for overcoming forgetting via forward pass. Based on these insights, we propose two easier and better enhancement to overcome forgetting and extend the application of related methods easily.
Impact Statement

### v4

@@ 4 ZeroFlow Benchmark
(c) FO-Adam
(d) ZO-Adam
(e) ZO-Adam ($q=4$)
(f) ZO-Adam-Sign
(g) ZO-Adam-Conserve
(h) FO-SGD
(i) ZO-SGD
(j) ZO-SGD ($q=4$)
(k) ZO-SGD-Sign
(l) ZO-SGD-Conserve
Figure 4: The Trajectory of Different Optimization during Overcoming Forgetting. , , and ★ denote the minima for the new, old, and both tasks, respectively. The trajectory is taken when using the total loss from both tasks (cyan).
@@
This section delves into the empirical performance of ZO optimization in overcoming catastrophic forgetting. Our ZeroFlow benchmark evaluates average performance across incremental stages, final-stage accuracy, forgetting, and efficiency, while accounting for dataset complexity and model diversity.
D.1 Benchmark Setups
@@
Forgetting scenarios, schemes, and models. We conduct evaluations under a standard catastrophic forgetting setting, namely class incremental learning. For this purpose, we investigate two state-of-the-art schemes: EASE and APER. Both models are initialized with ViT-B/16 pretrained on ImageNet-1K (IN1K), and are subsequently fine-tuned on four downstream tasks of varying complexity—ranging from standard benchmarks such as CIFAR-100 and CUB, to more challenging datasets like ImageNet-A and OmniBenchmark, which exhibit a large domain gap from the pretraining distribution (Zhou et al., 2024a, c). Following (Zhou et al., 2023a), each dataset is evenly split into 10 incremental tasks by class. For instance, OmniBenchmark contains 300 classes, with 30 classes introduced at each stage. No memory is permitted for storing past examples.
@@
Benchmark setup and details. To evaluate the application of ZeroFlow in forgetting scenarios, we include the methods described in Section C.1, specifically ZO (Ghadimi & Lan, 2013), Sign (Liu et al., 2019), and Conserve (Kim et al., 2021; Zhang et al., 2024), in comparison with their FO counterparts using SGD and Adam optimizers (Chen et al., 2019). Additionally, as highlighted in (Zhang et al., 2024), Forward-Grad (Baydin et al., 2022) which relies on forward mode automatic differentiation, potentially becomes a missing but competitive forward pass baseline. In a nutshell, ZeroFlow covers 7 forward pass-based methods: ZO-SGD, ZO-SGD-Sign, ZO-SGD-Conserve, ZO-Adam, ZO-Adam-Sign, ZO-Adam-Conserve, Forward-Grad. Unless otherwise specified, the query budget is fixed to 1 for efficiency. Notably, here we consider generating one set of perturbation vectors for the entire model as one query. In other words, we usually require 2 forward propagations for two-point finite difference gradient estimations.
@@
Evaluation metrics. Overall, we adopt two categories of evaluation metrics in ZeroFlow: accuracy and efficiency. The accuracy metrics include average accuracy across all tasks, final-task accuracy, and a forgetting score (BWT in Appendix B.5). The efficiency metrics encompass memory usage (GPU), query budget, and runtime. Together, these metrics provide insights into the resource demands of ZO optimization for mitigating forgetting.
D.2 Evaluation Results of ZeroFlow
ZeroFlow evaluation on continual learning. In Table 1, we evaluate the performance of different BP-free and BP-based (FO-SGD and FO-Adam) methods in a typical forgetting scenario (continual learning). We use two SOTA models as examples (EASE (Zhou et al., 2024b) and APER (Zhou et al., 2023a)) and investigate SGD and Adam optimizers, 7 forward pass-based methods, and four commonly used datasets. Several observations are listed below,
ZeroFlow helps manage memory and runtime. In Table 2, we compare the efficiency of various ZO and FO optimizers in mitigating catastrophic forgetting, focusing on two key aspects: memory cost (in GB) and runtime cost (in seconds). First, naive ZO optimization reduces memory usage by approximately fivefold compared to FO optimization. Moreover, ZO methods reduce runtime per iteration by around 50% relative to FO, significantly improving their practicality for overcoming forgetting. Notably, we regenerate the perturbation vectors for model parameters iteratively by storing random seeds. This degrades the vector granularity from full-model to per-layer level, thereby further reducing the memory required for forward evaluations in ZeroFlow, at the cost of additional runtime for regenerating the vectors. Second, the ZO and Sign variants demonstrate comparable efficiency in both memory and runtime. Although increasing the number of queries can impact runtime efficiency, it does not compromise memory advantages. Third, Conserve also demonstrates efficient memory management, although its runtime is approximately twice as long as that of naive ZO. This may partly explain its stronger performance in some scenarios, as shown in Table 1. Finally, the Forward Gradient method requires more memory than other ZO-based approaches because it involves computing gradients via the Jacobian-vector product (JVP), which necessitates storing all intermediate activations during the forward pass. For models like ViT, this includes large attention maps and other intermediate representations. In contrast, naive ZO methods only require two forward passes on perturbed inputs and avoid storing these intermediate values, resulting in much lower memory usage.
@@ D.2 Evaluation Results of ZeroFlow
Trade-off between performance and query number. As shown in Figure 5, we investigate the impact of query numbers on optimization performance, comparing SGD and Adam optimizers in the zeroth-order setting. Both optimizers demonstrate improved performance as query numbers increase across {1,2,4,8,16,32}, suggesting that additional function evaluations enable more accurate gradient estimation. The results suggest that in scenarios where function evaluation costs are manageable, higher query numbers can yield substantially better performance, with Adam being particularly effective at leveraging the additional gradient information for enhanced optimization outcomes.
@@
5 Insights and Discussions
Figure 7: Effectiveness of Hybrid ZO in Overcoming Forgetting. In Hybrid ZO, backward benefits from forward passes.
As shown in Figure 4, we visualized the optimization trajectories of both forward passes and backpropagation methods. Our analysis reveals several key insights:
Distinct trajectory characteristics of FO and ZO. FO approaches (Figure 4, 4) show smoother optimization paths due to their access to exact gradient information. In contrast, ZO methods demonstrate varying degrees of exploration behavior through trajectory jitter. This exploration pattern is particularly pronounced in ZO-Adam variants compared to ZO-SGD variants, indicating that the base optimizer choice significantly influences the exploration-exploitation trade-off during optimization.
Path characteristics in ZO optimization. Comparing base ZO methods with their $q=4$ counterparts (Figure 4 vs 4, Figure 4 vs 4), we observe that increasing query numbers leads to smoother trajectories, suggesting that more queries help provide more stable gradient estimates. The Sign variants (Figure 4, 4) demonstrate more pronounced oscillations in their trajectories, particularly visible in the ZO-Adam-Sign case. In contrast, the conservative variants (Figure 4, 4) maintain relatively stable paths that better balance between the old and new task minima.
6 New Enhancement to Mitigate Forgetting
Table 3: Effectiveness of Historical Estimation in Mitigating Forgetting. Proportion of 0% denotes that the plain optimizer ZO-SGD. Bold indicates the best performance.
Metrics Proportion
0% 20% 40% 60% 80%
Avg 57.87 58.90 58.76 58.34 57.83
Last 48.32 49.04 48.84 48.42 48.10
Fgt 11.08 11.79 11.78 11.60 11.57
(a) FO-SGD
(b) ZO-SGD
Table 4: Effectiveness of Sparsity-induced Estimation in Overcoming Forgetting. Proportion of 0% denotes the plain ZO-SGD. Bold indicates the best performance.
Ratio 0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Avg 57.87 59.17 59.46 59.29 59.39 59.45 59.26 59.39 59.38 59.47
Last 48.32 48.58 49.05 48.72 48.91 49.24 49.11 49.05 49.11 49.24
Fgt 11.08 12.65 12.17 12.76 12.53 12.37 12.36 12.54 12.46 12.33
Table 5: Ablation Studies on the Effectiveness of Combining Enhancements.
Optimizer Hybrid Historical Sparsity Avg Last
FO-SGD - - - 61.24 51.02
ZO-SGD - - - 57.87 48.32
61.40(+3.53) 51.34(+3.02)
58.90(+1.03) 49.04(+0.72)
59.47(+1.60) 49.24(+0.92)
62.07(+4.20) 51.94(+3.62)
Complementary Enhancements: The results in Table 5 demonstrate that the proposed enhancements are not mutually exclusive and can be effectively integrated. Specifically, FO training can substantially benefit from subsequent fine-tuning with hybrid ZO optimization, as illustrated in Figure 7. Notably, the inherent instability of ZO with large step fluctuations can sometimes facilitate escaping local minima and encourage broader exploration, which in turn benefits FO convergence. Furthermore, incorporating historical gradients and sparsity perturbations contributes to mitigating forgetting and stabilizing the optimization process.
7 Conclusion
This paper introduces ZeroFlow, a benchmark study that probes a series of forward pass-based methods for overcoming catastrophic forgetting. This work resorts to an easier way (no need for backpropagation and activation storage) to overcome forgetting. Concretely, our benchmarks include various forward pass-based methods, forgetting scenarios, and evaluation metrics. We also reveal the overlooked optimization principles for overcoming forgetting via forward passes. Based on these insights, we propose two easier and better enhancement to overcome forgetting and extend the application of related methods easily.
Impact Statement
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 62495063. This work was supported in part by the China Postdoctoral Science Foundation under Grant 2024M761677.
Appendix A Experimental Details
In this section, we provide an overview of zeroth-order optimization algorithms and the function settings used for the trajectory analysis.
A.1 Concise Overview of Zeroth-Order Estimation
Zeroth-order optimization aims to minimize/maximize an objective function $f:\mathbb{R}^{n}\to\mathbb{R}$ without derivative information. The core problem is formulated as $\min_{\theta\in\mathbb{R}^{n}}L(\theta)$, where $\theta$ denotes the optimization variable. To enable gradient-based updates, Simultaneous Perturbation Stochastic Approximation (SPSA(Spall, 1992)) is a commonly used technique to approximate gradients by perturbing the input variables. Specifically, the gradient ^$\hat{\nabla}L(\theta)$ at point $\theta$ is estimated as:
^$\hat{\nabla}L(\theta,\xi;B)=\frac{L(\theta+\epsilon\xi;B)-L(\theta-\epsilon\xi;B)}{2\epsilon}\cdot\xi^{-1},$ (4)
where $\xi\sim\mathcal{N}(\bm{0},\bm{I})$ is a random perturbation vector, and $\epsilon>0$ is a small perturbation step size (typically adjusted during optimization).
ZO-SGD(Ghadimi & Lan, 2013): Using the gradient estimator ^$\hat{\nabla}L(\theta,\xi;B)$, zeroth-order algorithms, such as ZO-SGD, follow the iterative update rule:
$\theta_{t+1}=\theta_{t}-\eta_{t}\cdot\hat{\nabla}L(\theta_{t},\xi_{t};B),$ (5)
where $\eta_{t}$ is the learning rate at step $t$. ZO-SGD bypasses explicit gradient computation through local function evaluations, making it suitable for high-dimensional, non-convex optimization problems.
ZO-SGD-Sign(Liu et al., 2019): A variant of ZO-SGD, known as ZO-SGD-Sign, improves upon the original approach by approximating the gradient direction using the sign of the gradient estimate. The update rule becomes:
$\theta_{t+1}=\theta_{t}-\eta_{t}\cdot\text{sign}(\hat{\nabla}L(\theta_{t},\xi_{t};B)),$ (6)
where $\text{sign}(\cdot)$ denotes the element-wise sign function. This approach often leads to faster convergence in some problems where the magnitude of the gradient is not as important as its direction.
ZO-SGD-Conserve(Bergou et al., 2020): ZO-SGD-Conserve is another variant that conservatively selects the update direction by locally comparing three candidate points, rather than directly committing to a single gradient step. The update rule for this method is:
$\theta_{t+1}=\arg\min_{y\in\mathcal{C}_{t}}f(y),\quad\mathcal{C}_{t}=\left\{\theta_{t},\ \theta_{t}-\eta_{t}\cdot\hat{\nabla}L(\theta_{t},\xi_{t};B),\ \theta_{t}+\eta_{t}\cdot\hat{\nabla}L(\theta_{t},\xi_{t};B)\right\},$ (7)
This method mitigates overly aggressive updates by evaluating possible directions and choosing the one that locally minimizes the objective function.
ZO-Adam(Zhang et al., 2024): ZO-AdaMM (Chen et al., 2019) is the first attempt to apply the Adam family (specifically AMSGrad(Reddi et al., 2019)) to zeroth-order (ZO) optimization algorithms, providing convergence guarantees for both convex and nonconvex settings. The update rule is given by:
$\theta_{t+1}=\theta_{t}-\eta_{t}\cdot\frac{m_{t}}{\sqrt{V_{t}}+\epsilon},\quad V_{t}=\mathrm{Diag}(\max(v_{t},v_{t-1})),$ (8)
$m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\hat{\nabla}L(\theta_{t},\xi_{t};B),\quad v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\,(\hat{\nabla}L(\theta_{t},\xi_{t};B))^{2},$
In our implementation, we simply replace SGD with Adam for convenience, referring to this variant as ZO-Adam. Nevertheless, we also provide a reference implementation of the original oracle ZO-AdaMM algorithm.
Forward Gradient Descent (FGD)(Baydin et al., 2022): FGD replaces backpropagation with forward-mode automatic differentiation to estimate gradient directions using Jacobian-vector products (JVPs). Instead of computing full gradients via reverse-mode automatic differentiation (AD), FGD samples probe vectors to construct unbiased estimators of the gradient direction. A typical FGD update step is:
$\theta_{t+1}=\theta_{t}-\eta_{t}\cdot\text{JVP}_{\theta_{t}}(v_{t})=\theta_{t}-\eta_{t}\cdot\left.\frac{df(\theta)}{d\theta}\cdot v_{t}\right|_{\theta=\theta_{t}},$ (9)
where $v_{t}$ is a random probe vector (e.g., Rademacher or Gaussian), and $\text{JVP}_{\theta_{t}}(v_{t})$ represents the forward-mode gradient approximation in direction $v_{t}$. FGD enables training when reverse-mode AD is impractical or unavailable, and offers flexibility for hardware or software systems that only support forward execution. We denote Forward as FGD throughout this paper.
A.2 Function Settings
Following the setup in CAGrad (Liu et al., 2021), we visualize the optimization behavior of first-order (FO) and zeroth-order (ZO) methods in mitigating forgetting. Specifically, we consider a two-dimensional parameter space $\theta=(\theta_{1},\theta_{2})\in\mathbb{R}^{2}$, with the following task-specific loss functions: $L_{1}(\theta)=c_{1}(\theta)f_{1}(\theta)+c_{2}(\theta)g_{1}(\theta)$ for the old task (orange), and $L_{2}(\theta)=c_{1}(\theta)f_{2}(\theta)+c_{2}(\theta)g_{2}(\theta)$ for the new task (red). The parameter point is initialized at $[-8.5,-5]$ to be closer to old tasks, facilitating better adaptation to them. The contour plot in Figure 3 illustrates the overall objective function defined as $L(\theta)=L_{1}(\theta)+L_{2}(\theta)$, where the $x$- and $y$-axes correspond to $\theta_{1}$ and $\theta_{2}$, respectively.
[…]
$g_{1}(\theta)=\frac{(-\theta_{1}+7)^{2}+0.1(\theta_{2}-8)^{2}}{10}-20,$
$g_{2}(\theta)=\frac{(-\theta_{1}-7)^{2}+0.1(\theta_{2}-8)^{2}}{10}-20,$
[…]
Appendix B Additional Results
B.1 Comprehensive Analysis of Memory Usage on ZeroFlow
Figure 9: Comparison of Memory Usage between FO-SGD and ZO-SGD with Different Batch Sizes. $\Delta$ denotes the increase in memory usage of the final task compared to the initial one.
In this subsection, we provide a detailed comparison of memory usage during incremental learning to demonstrate the storage efficiency of ZeroFlow (ZO-SGD) compared to its counterpart, FO-SGD. Figure 9 illustrates the peak memory usage of MEMO when trained on the CIFAR-100 dataset. The backbone is fixed as a pretrained ViT-B/16-IN1K model, which is subsequently fine-tuned with batch sizes ranging from 64 to 512. The experimental results highlight the following observations:
First, doubling the training batch size significantly increases the memory consumption of FO-SGD, requiring more GPU resources. For instance, completing the entire incremental training process on FO requires 1, 2, 3, and 6 GPUs, respectively, for batch sizes of 64, 128, 256, and 512, with each GPU equipping with 24GB of memory. In contrast, ZO-SGD training consistently requires only one GPU resource.
Second, as training progresses, the memory demands for larger batch sizes increase rapidly. For FO, the memory consumption for 512 batches at stage 5 grows by 30.08 GB compared to the initial stage. In contrast, ZO-SGD shows a modest increase of only 3.92 GB, maintaining a low growth rate. As training advances, the memory efficiency of ZO-SGD becomes more pronounced, especially for model-expansion based CL models.
B.2 More Observations on Optimization Trajectories during Overcoming Forgetting
@@
(k) FO-Adam
(l) ZO-Adam
(m) ZO-Adam ($q=4$)
(n) ZO-Adam-Sign
(o) ZO-Adam-Conserve
(p) FO-SGD
(q) ZO-SGD
(r) ZO-SGD ($q=4$)
(s) ZO-SGD-Sign
(t) ZO-SGD-Conserve
@@
Figure 10: The Trajectory of Different Optimization during Overcoming Forgetting. The first and last two rows are trained for 100k steps with learning rates of 0.001 and 0.01, respectively. Red denotes the minimum of new task, orange denotes the minimum of old task. The cyan trajectory taken when using the total loss from both tasks.
In this subsection, we present a different scenario where the model is initialized at a local minimum $\theta_{1},\theta_{2}=\{-4.0,5.0\}$, surrounded by intricate valleys, but training with different learning rate as shown in Figure 10. For a learning rate of 0.001, the first-row subfigures demonstrate that Adam using both FO and ZeroFlow stagnate in the valley. Even with bias correction, the Adam optimizer still fails to escape the local region without sufficient momentum. However, ZO-Adam-Sign seems to successfully optimize towards the region around the global minimum. Unlike ZO-Adam, ZO-Adam-Sign applies the gradient using a sign function, which outputs either +1 or -1 depending on the gradient direction. This discrete update method, which lacks continuous gradient information, causes ZO-Adam-Sign to take larger, step-like jumps. Particularly in the early stages, where gradient information is sparse or noisy, this leads to more fluctuations and introduces greater randomness in the optimization process, helping it to cross over the valleys. The second-row subfigures use SGD as the base optimizer. We observe that, except for ZO-SGD-Sign, both ZeroFlow and FO-SGD converge effectively. This can be attributed to SGD’s simple update rule based on function values. Notably, FO-SGD escapes the valley by leaping to a higher and flatter region, while ZeroFlow demonstrates the ability to traverse beneath valleys. With a higher learning rate of 0.01, FO-Adam, ZO-Adam with four queries, and ZO-Adam-Sign escape the local region more easily. However, ZO-Adam still stagnates along the valley, demonstrating the stabilizing effect of multiple query loops. Similarly, ZO-Adam-Conserve suffers from the risk of an overly conservative strategy. ZO-SGD also fails to converge to the optimum due to gradient explosion caused by the large learning rate. In contrast, ZeroFlow shows minimal degradation despite its inherent randomness.
As a result, the behavior of ZeroFlow—sometimes escaping the valley but failing to converge to the optimum, and sometimes getting trapped with low query counts but not with higher ones—highlights the trade-off between randomness and stability during updates. With larger search loops, lower learning rates, and more stable update steps, the model becomes increasingly prone to getting stuck in local minima, especially in continual learning scenarios where balancing old and new tasks introduces additional complexity.
B.3 Extra Evaluation on Memory Replay Methods
We further evaluate the performance of ZeroFlow when applied to a representative replay-based method (MEMO (Zhou et al., 2023c), replay buffer = 2000), to demonstrate its broader applicability. As shown below, ZeroFlow consistently remains stable in mitigating forgetting. Notably, although the average accuracies exhibit slight gaps compared to FO methods, the accuracies at the final stage progressively approach or even surpass those of the FO baselines on the CIFAR-100 dataset.
Table 6: Accuracy Results on MEMO.
Method Optimizer Strategy CIFAR-100 ImageNet-A
Avg Last Avg Last
MEMO SGD FO 87.43 79.66 53.15 38.97
ZO 85.92 79.00 46.87 25.81
Sign 85.72 79.10 53.31 38.18
Conserve 85.86 79.20 47.20 28.51
Adam FO 86.45 76.17 54.06 41.54
ZO 85.86 78.59 52.70 39.01
Sign 86.16 76.38 53.10 39.82
Conserve 85.89 77.71 53.20 39.57
- Forward 84.63 76.32 53.59 40.64
B.4 Memory and Time Efficiency on Larger Transformers
To assess the scalability of ZeroFlow, we evaluated its efficiency on two larger vision transformers, ViT-L/16 and ViT-H/14. As shown below, ZeroFlow consistently offers substantial memory savings across all model sizes. Notably, even when using ZO-SGD-Sign, the runtime remains faster than that of standard FO optimization.
Table 7: Evaluation on lager transformers.
Optimizer ViT-B/16 ViT-L/16 ViT-H/14
[…]
FO-SGD 12.08GB 59.3s 33.27GB 65.0s 78.09GB 190.1s
ZO-SGD (q=1) 2.41GB 32.4s 3.77GB 47.0s 6.45GB 118.7s
ZO-SGD (q=4) 2.41GB 111.7s 3.77GB 178.3s 6.45GB 442.6s
ZO-SGD-Sign 2.41GB 32.4s 3.77GB 48.7s 6.45GB 119.3s
ZO-SGD-Conserve 2.41GB 70.1s 3.77GB 108.9s 6.45GB 222.3s
Forward 3.94GB 45.9s 5.82GB 142.0s 9.85GB 372.5s
B.5 Longer Task Sequence
To further assess robustness, we evaluate performance on an extended task sequence consisting of 20 tasks. As shown below, ZeroFlow continue to deliver comparable performance. Additionally, following (Wang et al., 2024), we additionally adopt the FWT and BWT metrics to assess the overall performance of ZeroFlow. FWT (Forward Transfer) quantifies the average influence of prior knowledge on the learning of new tasks, while BWT (Backward Transfer) measures the average influence of learning new tasks on the performance of previously learned $K-1$ tasks.
$\text{FWT}=\frac{1}{K-1}\sum_{j=2}^{K}(a_{j,j}-\tilde{a}_{j}),\quad\text{BWT}=\frac{1}{K-1}\sum_{j=1}^{K-1}(a_{K,j}-a_{j,j})$ (10)
Here, $a_{k,j}$ denotes the accuracy on task $j$ after training on the $k$-th dataset, and $\tilde{a}_{j}$ represents the accuracy of a random initialized model trained only on dataset $\mathbb{D}_{j}$.
Table 8: Additional Experimental Results of EASE on 20 Sequential Tasks.
Method Optimizer Strategy Avg Last FWT BWT
EASE SGD FO 87.32 80.20 -6.89 -6.79
ZO 82.65 75.98 -8.33 -7.71
Sign 83.47 76.13 -8.01 -7.22
Conserve 82.20 75.94 -8.64 -7.93
Adam FO 86.67 78.19 -7.17 -6.80
ZO 84.07 76.89 -7.92 -7.19
Sign 84.16 76.90 -7.95 -7.20
Conserve 83.82 76.76 -8.04 -7.07
- Forward 82.84 76.32 -8.25 -10.84